At an income of 100,000 a consumer spent 90 000 on consumer goods.when income rose to 200,000 the consumer spent 160 000 on consumer goods what is the MPC

1
Expert's answer
2022-01-30T14:07:19-0500

Answer


MPC= ΔC / ΔY

ΔC is the change in consumption

ΔY is the change in income.

ΔC=160,000-90,000

=70,000

ΔY=200,000-100,000

=100,000

"MPC=\frac{70,000}{100,000}=0.7"